1.)   A solution of 0.40 M HCl is found to have a pH of 0.52. Know that pH, when properly defined, is given in terms of activities, or pH º - log10[aH+].

a.)    Calculate the ideal solution pH, the activity of this solution and the activity coefficient.

b.)   Use the Debye-Huckel Limiting Law to calculate the pH and compare to experimental and ideal. Explain any differences.
Answer #1

HCl + H2O <==> H3O+ + Cl-

pH = -log(aH+)

In terms of activity we have to:

0.52 = -log(aH+)

100.52 = 10-log(aH+)

aH+ = 3.31 M

--

From activity we can find:

aH+ = xh+ * yh+

where:

xh+ is the mole froction of the mixture

yh+ is the activity coefficient

yh+ = 3.31M/0.40M

yh+ = 8.275

---

Ideal pH

HA <==> H+ + A-

0.40 0 0

-0.40 +0.40 +0.40

0.00 0.40 0.40

[H+] = 0.40 M

pH = -log(H+) = -log(0.40)

pH = 0.39

---

b) The Duckel-Huckel limiting law

The Debye-Hückel Limiting Law, is applicable only for very low concentrations of solute (until unless contributions of the non-ideality become important). This law helps us in the calculation of the mean activity coefficient from basic properties of the solution:

Log γi = -AZi2√I

where;

z+ and z- are the charge number of respectively the cation and the anion concerned.

A is an empirical parameter, dependant upon the solvent and temperature (eg, for a solution in water at 298K, A = 0.509).

I is the ionic strength of the solution,

I =1/2 [\sumCiZi2]

Cl- charge = -1

H+ charge = +1

I = 1/2 (Concentration of HCl × # of H ions × (Charge of H)2) + (Concentration of HCl in M × # of Chloride ions × (Charge of Chloride)2)]

I = 1/2 [(0.40 M × 1 × (+1)2) + (0.40 M × 1 × (−1)2)] = 0.80 M / 2 = 0.40 M

ln yH+ = -0.509 * |(+1)(-1)| * sqrt(0.40)

eln yH+ = e0.322

yH+ = 1.37

aH+ = 0.40 * 1.37

aH+ = 0.552

pH = -log(aH+)

pH = -log(0.552)

pH = 0.25